A circular loop of wire having a radius of 5.94 cm carries a current of 0.280 A. A vector of unit length and parallel to the dipole moment of the loop is given by 0.60i – 0.80j . If the loop is located in a uniform magnetic field given by Upper B = (0.715 T) i + (0.731 T) k find (a) the x-component, (b) the y-component, and (c) the z-component of the torque on the loop and (d) the magnetic potential energy of the loop.
Answer #1

\tau = \mu \ \textup{x} \ B

  =I A \mu^{'} \textup{x} \ B

To calculate the cross product, using the determinant method

i x j = k, j x k = i, k x i = j, a x b = - b x a, and a × a = 0 unless both vectors have all nonzero components.

\tau =I A \mu^{'} \textup{x} \ B

= IA [ (0.60i - 0.80j) × (0.715 i + 0.731 k)]

= IA [ -0.5848 i -0.4386j +0.572k ]

Using values I = 0.28 and A = pi * 0.0594^2 = 0.011 m^2

Torque = (-1.8011 * 10^-3 ) i - (1.3508 * 10^-3 )j +(1.7 * 10^-3) k N/m